The average number of automobiles per minute stopping for gas at a particular service station along the Coastal Road is 4. What is the probability that in any given minute more than two will stop for gas?

1
Expert's answer
2022-04-19T05:07:31-0400

It has to be pointed out that the distribution is not specified. Suppose that "X" is a random variable that corresponds to the number of automobiles per minute. The aim is to compute "P(X>2)". Consider the most typical distributions:

  1. Suppose that "X" has a normal distribution with parameters "\mu=4" and "\sigma" . Then, "P(X>2)=\int_{2}^{+\infty}f(x)dx", where "f(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12\left(\frac{x-\mu}{\sigma}\right)^2}". For "\sigma=1" we get: "P(X>2)\approx0.977."
  2. Suppose that "X" has a Poisson distribution with "\lambda=4." Remind that for Poisson distribution "P(X=k)=\frac{\lambda^ke^{-\lambda}}{k!}". Thus, "P(X>2)=P(X=3)+P(X=4)+...\approx0.762"

Thus, for the normal distribution we receive "0.977" and for the Poisson distribution "0.762".